• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 08:01:13 on June10,2010

    Entry number 325531

    keyword=Owl shift summary 06/10/10

    Shift summary: Owl, 06/10/2010

    00:00 - Shift starts. Right arm down, middle of run 4443. Note that prompt output does not accurately show beam current or # of mC.
    00:15 - Checking HVs. See the right arm scintillators (channels 9.5, 9.6) on. Turned off.
    00:20 - See dithering off. Wait five minutes for run 4443 to end, then turn it on.
    04:40 - Compton rate flatlined ~15 minutes ago. Followed the Compton ops manual, rebooted ROC6. Seems fine now. Will stop/start new run to re-synchronize it with the parity daq.
    07:45 - Cryo group called to discuss the problems with the Moller magnet. Don't know much about the magnet status, referred him to Jixie.
    08:00 - Passing off to next shift.

    Runs (all have dithering and feedback on, unless otherwise noted):
    4443 - Started on swing, no dithering, yes feedback. 66.9k multiplets
    4444 - 78.3k multiplets
    4445 - 58.9k multiplets
    4446 - 69.6k multiplets
    4447 - 59k multiplets
    4448 - 68.8k multiplets
    4449 - 63k multiplets
    4450 - 72.5k multiplets
    4451 - 79 multiplets
    4452 - In progress at the end of shift.

    Shift totals: 616k multiplets, 514 mC (approximated by # of multiplets @ 50uA, accounting for factor of 1/2 for L. Arm only)